OpenStudy (anonymous):

Solve cos 2 θ = cos θ + 2 for 0 < θ < 2p

OpenStudy (anonymous):

@ikram002p what would we do in this situation

OpenStudy (ikram002p):

is it \(\Huge \cos^2 \theta \) or \(\Huge \cos 2 \theta \)

OpenStudy (anonymous):

it is cos 2θ

OpenStudy (anonymous):

so could we divide cos θ on each side and it would be θ=2

OpenStudy (anonymous):

@ikram002p what happens

OpenStudy (ikram002p):

sry my account acting weird :-|

OpenStudy (ikram002p):

so lets do it cos 2θ = 2 cos^2 θ -1 right ?

OpenStudy (ikram002p):

cos 2 θ = cos θ + 2 2 cos^2 θ -1 = cos θ + 2

OpenStudy (anonymous):

no Solve cos 2 θ = cos θ + 2 for 0 < θ < 2p

OpenStudy (anonymous):

ya

OpenStudy (ikram002p):

i know i know but cos 2 θ = 2 cos^2 θ -1

OpenStudy (ikram002p):

did u got this one ?

OpenStudy (anonymous):

no i don't

OpenStudy (ikram002p):

shall i keep going ?

OpenStudy (ikram002p):

ok ill tell you what it is

OpenStudy (anonymous):

why does it become from cos 2 θ = cos θ + 2 to cos 2 θ = 2 cos^2 θ -1

OpenStudy (ikram002p):

do u know this rol cos^2 + sin^2=1 ??

OpenStudy (anonymous):

no

OpenStudy (ikram002p):

ohh u dnt know it :-\

OpenStudy (ikram002p):

ok u know how to solve cos (A+B ) ??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so from cos 2 θ = cos θ+2 what will we do next

OpenStudy (ikram002p):

ughh no !

OpenStudy (ikram002p):

cos 2 θ = cos ( θ+θ) right ??

OpenStudy (anonymous):

ya

OpenStudy (ikram002p):

cos ( θ+θ) = cos θ cos θ - sin θ sin θ right ?

OpenStudy (anonymous):

yup

OpenStudy (ikram002p):

cos 2 θ = cos ( θ+θ)=cos^2 θ - sin^2 θ u agree with this untill nw ?

OpenStudy (anonymous):

ya

OpenStudy (ikram002p):

ok now use this law ( u have to memorise it ) cos^2 θ + sin^2 θ=1 ok ?

OpenStudy (anonymous):

can we subtract cos θ from one side to the other so the equation can be become cos θ=2

OpenStudy (anonymous):

then θ=cos^-1(2)

OpenStudy (ikram002p):

no u cant coz u have 2thata in one side and theta in the sec side

OpenStudy (ikram002p):

u agree to that ?

OpenStudy (anonymous):

ya

OpenStudy (ikram002p):

u only can remove it when both angels are the same ok ?